In an azeotropic solution, the liquid and vapor phases have equal mole fractions of solute at equilibrium. Find the numerical value of the Henrt's Law coefficient that is azeotropic at its standard boiling point. Include units.
Answer #1

50% Vapoo 1 phase liquid Temp 2 phase liquid XB → Two phases of liquid Creating a minimum azotropic mature “The azeotropic miBenzene a water Mole percent system. of water = 29.6% Mole of. of Benzene - - 100-29.6 - 70.4 % = Xo Mole Dercent of SolutionXp = 89.4/100 = 0.694 13= km (0.894) kn = 111x105 Thus, henry Constant value is lol1 x 105.Azeotropic Solution : → A mixture of Liquids a constant boiling point. that has has Because, an azeotropic miature may be higThe Solution mixture formed from Dositive deviation Desults in minimum boiling azeotrope Ethanol - Wateo at 89.4 mole Peocent